# SemVideo: Reconstructs What You Watch from Brain Activity via Hierarchical Semantic Guidance

### v1

@@ 4.1 Main Results
Following the work of [43], we reconstruct video clips using Tune-a-video [77] which lacks motion priors and conduct shuffle tests to further validate the contributions of both the $MAD$ and the motion-oriented narratives $C_{motion}$ to motion improvement. Specifically, for each 12-frame reconstructed video, we randomly shuffle the order of the frames 100 times and compute the EPE on both the original and shuffled frames. We estimate the P-value using the following formula: $P=\sum_{i=1}^{100}\delta_{i}/100$, where $\delta_{i}=1$ if the $i^{th}$ shuffle outperforms the reconstruction result in the original order based on the metrics, and $\delta_{i}=0$ otherwise. A lower P-value signifies a closer alignment between the sequential order of the reconstructed video and the ground truth. It can be observed that the P-value for our full model is significantly low, confirming that its predicted inter-frame order robustly matches the ground truth, but rises substantially when $C_{motion}$ is removed and spikes to 0.84 when the entire $MAD$ is ablated, as illustrated in Figure.5 left. This indicates that we have indeed decoded motion information from $MAD$ and fMRI-derived $C_{motion}$. This conclusion is reinforced by a comparison of the EPE score, as shown in Figure.5 right. The EPE increase when removing $C_{motion}$ and the spike when ablating the $MAD$ confirm the crucial role of motion semantic guidance and that the $MAD$, not T2V priors, is the source of decoded motion improvement.
@@ D.2 Failure cases
To facilitate a comprehensive and objective assessment of $SemVideo$, we analyze several of its reconstruction failure cases in Figure.11. A primary cause of these failures stems from the inherent data acquisition paradigm. Specifically, the experimental data is created by evenly segmenting longer videos, which introduces abrupt content transitions at the boundaries of the resulting clips. Our model, $SemVideo$, struggles to replicate these sudden shifts in the reconstructed outputs.

### v2

# arXiv:2602.21819v2[cs.CV] 27 Feb 2026

@@ 4.1 Main Results
Following the work of [43], we reconstruct video clips using Tune-a-video [77] which lacks motion priors and conduct shuffle tests to further validate the contributions of both the MADMAD and the motion-oriented narratives $C_{motion}$ to motion improvement. Specifically, for each 12-frame reconstructed video, we randomly shuffle the order of the frames 100 times and compute the EPE on both the original and shuffled frames. We estimate the P-value using the following formula: $P=\sum_{i=1}^{100}\delta_{i}/100$, where $\delta_{i}=1$ if the $i^{th}$ shuffle outperforms the reconstruction result in the original order based on the metrics, and $\delta_{i}=0$ otherwise. A lower P-value signifies a closer alignment between the sequential order of the reconstructed video and the ground truth. It can be observed that the P-value for our full model is significantly low, confirming that its predicted inter-frame order robustly matches the ground truth, but rises substantially when $C_{motion}$ is removed and spikes to 0.84 when the entire MADMAD is ablated, as illustrated in Figure.5 left. This indicates that we have indeed decoded motion information from MADMAD and fMRI-derived $C_{motion}$. This conclusion is reinforced by a comparison of the EPE score, as shown in Figure.5 right. The EPE increase when removing $C_{motion}$ and the spike when ablating the MADMAD confirm the crucial role of motion semantic guidance and that the MADMAD, not T2V priors, is the source of decoded motion improvement.
@@ D.2 Failure cases
To facilitate a comprehensive and objective assessment of SemVideoSemVideo, we analyze several of its reconstruction failure cases in Figure.11. A primary cause of these failures stems from the inherent data acquisition paradigm. Specifically, the experimental data is created by evenly segmenting longer videos, which introduces abrupt content transitions at the boundaries of the resulting clips. Our model, SemVideoSemVideo, struggles to replicate these sudden shifts in the reconstructed outputs.